Circular No. 8426

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 V1 (SKIFF) B. Skiff, Lowell Observatory, reports his discovery of a comet on LO-

NEOS images obtained on Nov. 4.1 UT (discovery observation given below) with the 0.59-m Schmidt telescope, the object showing a moderately condensed coma of diameter 25′′ and a weak tail 50′′ long in p.a. 75^{o}. Following posting on the “NEO Confirmation Page”, B. L. Stevens (Las Cruces, NM, 0.3-m Schmidt-Cassegrain telescope) reports that his CCD images taken on Nov. 4.2 show a 30′′ tail in p.a. 55^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Nov. 4.06749  21^{h}08^{m}06.^{s}21  −0^{o}34′05′′.5  17.9

The available astrometry (including prediscovery LINEAR observations from Oct. 7 and 18), the following preliminary orbital elements, and an ephemeris appear on MPEC 2004-V16.

T = 2004 Dec. 8.861 TT  ω = 144.^{o}765
e = 0.69415  Ω = 242.236 }2000.0
q = 1.41841 AU  i = 11.467

a = 4.63754 AU n^{o} = 0.098690 P = 9.99 years

COMETS C/2004 S2, C/2004 S3, C/2004 S4 (SOHO) Additional Kreutz sungrazers (continuation to IAUC 8419 and 8423;

YT = Y.-s. Tsai):

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 S2 Sept.20.356 11^{h}27.^{m}1 − 0^{o}07′ C3  KB 2004-V05
C/2004 S3  22.959 11 37.5 − 1 08 C3/2 YT 2004-V05
C/2004 S4  28.529 12 03.0 − 3 16 C3/2 HS 2004-V05

SUPERNOVA 2004fe IN NGC 132 M. Modjaz, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrogram (range 390–900 nm) of SN 2004fe (cf. IAUC 8425), obtained by P. Challis, A. Garg, and C. Stubbs on Nov. 3.02 UT with the Magellan Clay 6.5-m Telescope (+ LDSS-2), shows it to be a supernova of type Ic. The spectrum is similar to spectra of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11) taken a few days before maximum.

2004 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green